The Evanec Companys next expected dividend, D1, is $2.91; its growth rate is 6%; and its common stock now sells for $35.00.
Answer #1

a.Cost of retained earnings=(D1/Current price)+Growth rate

=(2.91/35)+0.06

=14.31%(Approx)

b.% flotation cost=(Current price-Net price)/Current price

=(35-31.5)/35

=10%

c.Cost of new common stock=(D1/Net price)+Growth rate

=(2.91/31.5)+0.06

=15.24%(Approx)